### v1

@@ 3.3.1 Regularization-based approach
1:Input: $x^{0}$, $\rho_{I}>0$, $\kappa\in(0,1)$, $\tau^{0}>0$, $N^{\tau}>1$For $k=0,\ldots,N^{\tau}$
@@
[…]
5:Solve $\mathrm{LPEC}(x^{*}(\tau^{k}),\rho_{I})$ for $d^{k}$If $d^{k}$ exists
[…]
A regularization-based Phase I is summarized in Algorithm 2. In a practical implementation, a point is declared feasible if the constraint residual is below a certain tolerance. Formally, we require $h(x)\leq\texttt{tol}_{h}$, where $h(x)$ is the total infeasibility defined as:
$h(x)=\|(h_{c}(x),h_{\perp}(x))\|_{\infty},\ h_{c}(x)=\|c^{-}(x)\|_{\infty},\ h_{\perp}(x)=\max(|\min(x_{1},x_{2})|),$
with $c^{-}(x)=\min(0,c(x))$.
In our implementation, we take a finite number $N^{\tau}$ of homotopy steps. If the number of steps is large enough, Algorithm 2 eventually finds a feasible point or certifies local infeasibility. If $\mathrm{Reg}(\tau^{k})$ is infeasible, or its solution is already feasible for the MPEC, the algorithm terminates. Otherwise, the solution $x^{*}(\tau^{k})$ is used to construct ${\mathrm{LPEC}}(x^{*}(\tau^{k}),\rho_{I})$. The trust region radius $\rho_{I}$ must be sufficiently large to ensure that this LPEC is feasible. Note that a solution of $\mathrm{Reg}(\tau^{k})$ already satisfies the general inequality constraints, i.e., $c(x^{*}(\tau^{k}))\geq 0$. Therefore, we focus on values that make the complementarity constraint in the LPEC feasible. In our implementation, we use a fixed trust region radius with $\rho_{I}>0$. The smallest value of $\rho_{I}$ needed to make the LPEC feasible is $\rho_{I}>h_{\perp}(x^{*}(\tau^{k}))$, and we attempt to solve the LPEC only if this inequality is satisfied. If the LPEC is infeasible, the homotopy parameter is reduced, to obtain a better linearization point $x^{*}(\tau^{k})$. If a solution to the LPEC is found, Algorithm 2 attempts to solve the corresponding BNLP. If this is successful, a feasible point is found, otherwise, the homotopy parameter is reduced to obtain a better linearization point $x^{*}(\tau^{k+1})$. We prove in Theorem 4.1 that if $x^{*}(\tau^{k})$ is sufficiency close to a feasible point $x\in\Omega$, then the $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{*}(\tau^{k})+d^{k}),{\mathcal{I}}_{2}(x^{*}(\tau^{k})+d^{k}))$ is a feasible BNLP, where $d^{k}$ is a feasible point of $\mathrm{LPEC}(x^{*}(\tau^{k}),\rho^{I})$ In our implementation, every $\mathrm{Reg}(\tau^{k})$, for $k>0$ is initialized with the solution of $\mathrm{Reg}(\tau^{k-1})$, and $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{*}(\tau^{k})+d),{\mathcal{I}}_{2}(x^{*}(\tau^{k})+d))$ is initialized with $x^{*}(\tau^{k})$. The constants in our implementation are $\kappa=0.1$, $\tau^{0}=1$, $N^{\tau}=12$, ${\rho}^{I}=10^{-1}$, and $\texttt{tol}_{h}=10^{-10}$.
Solving the LPECs in Algorithm 2 can be computationally expensive. Thus, we may look for other strategies for guessing the active sets ${\mathcal{I}}_{0+}^{0}$, ${\mathcal{I}}_{0+}^{0}$, and ${\mathcal{I}}_{00}^{0}$ corresponding to a feasible point $x^{0}$ which is close to $x^{*}(\tau^{k})$. Some of such strategies are studied in Lin2006 . They all resemble the active set guessing in (19), but where $\epsilon$ is replaced by some function of $x$ and $\tau$. These strategies can be effective if the set ${\mathcal{I}}_{00}^{0}$ is an empty set. Otherwise, strong assumptions are needed for a correct active set identification if ${\mathcal{I}}_{00}\neq\emptyset$. Unfortunately, slightly over or underestimating ${\mathcal{I}}_{00}^{0}$ can quickly lead to an infeasible BNLP. For example, taking ${\mathcal{I}}_{00}=\{1\}$ in Example 3.4, and doing a wrong partition leads to an infeasible BNLP.
For the sake of comparison, we implement a simple LPEC-free strategy, which replaces lines 4-5 in Algorithm 2 with the following active set identification strategy. We use  (19) and set $\epsilon=\tau$. The biactive set is split into ${\mathcal{D}}_{1}(x(\tau))=\{i\in{\mathcal{I}}_{00}^{\tau}\mid x^{*}_{1,i}(\tau)<x^{*}_{2,i}(\tau)\}$ and ${\mathcal{D}}_{2}(x(\tau))=\{i\in{\mathcal{I}}_{00}^{\tau}\mid x^{*}_{1,i}(\tau)\geq x^{*}_{2,i}(\tau)\}$, which finally results in the following partition for the BNLP:
${\mathcal{I}}_{1}=\{i\mid x^{*}_{1,i}(\tau)<x^{*}_{2,i}(\tau)\},\quad{\mathcal{I}}_{2}=\{i\mid x^{*}_{1,i}(\tau)\geq x^{*}_{2,i}(\tau)\}.$ (24)
Note that in the overall algorithm, LPECs are still needed in Phase II.
3.3.2 Feasibility-based approach
A common approach for finding a feasible point in many optimization algorithms is to solve an auxiliary problem that minimizes the total infeasibility Fletcher2002 ; Waechter2006 .
We propose to solve the following feasibility problem:
$\underset{x\in{\mathbb{R}}^{n}}{\mathrm{min}}\;\quad$ $\|c^{-}(x)\|$ (25a)
s.t. $0\leq x_{1}\perp x_{2}\geq 0,$ (25b)
where $c^{-}(x)=\min(0,c(x))$ is the component-wise infeasibility of the inequality constraints. A crucial property of this MPEC is that it is always feasible, and we can directly apply Phase II of MPECopt to it. However, its objective is nondifferentiable. We discuss several special cases (in which we specify the norm in (25)), where this problem can be made smooth and solved efficiently in practice.
First, we can replace the objective of  (25) with $\|c^{-}(x)\|_{2}^{2}$, which is now once continuously differentiable. This allows us to apply a sequential LPEC method for MPECs with bound constraints Kirches2022 . Remarkably, the LPEC, in this case, has only bound and complementarity constraints, and it can be solved to global optimality with linear complexity (Kirches2022, , Proposition 2.4). However, the convergence rate of sequential LPEC methods is usually slow, and we may want to use Phase II of MPECopt directly. If there are only equality constraints but no inequality constraints in (1), then the squared two norm objective is twice continuously differentiable, and one can use Phase II of MPECopt combined with the fast explicit LPEC solver from (Kirches2022, , Proposition 2.4). In the presence of inequality constraints, one has to solve the usual LPEC as in Eq. (13).
Second, if we use the $\ell_{1}$-norm, a well-known smooth reformulation is:
$\underset{x\in{\mathbb{R}}^{n},s\in{\mathbb{R}}^{n_{c}}}{\mathrm{min}}\;\quad$ $\sum_{i=1}^{n_{c}}s_{i}$ (26a)
s.t. $c_{i}(x)\geq-s_{i},$ $\forall i\in\{1,\ldots,n_{c}\}$ (26b)
$0\leq x_{1}\perp x_{2}\geq 0,$ (26c)
$s\geq-s^{\mathrm{lb}}\,e,$ (26d)
where $e\in{\mathbb{R}}^{n_{c}}$ is a vector of all ones and $s^{\mathrm{lb}}$ is a nonnegative constant set here to zero. The $\ell_{\infty}$ norm is obtained similarly if we use a scalar slack variable instead of $s\in{\mathbb{R}}^{n_{c}}$. If we wish to find a strictly feasible point for the general inequality constraints, we can set $s^{\mathrm{lb}}$ to some positive number, and terminate the algorithm as soon as all components of $s$ are below zero. Since the MPEC (26) is always feasible we treat it with Phase II of MPECopt, outlined in the next section. If Phase II terminates with a B-stationary point and some $s_{i}>0$, we conclude that the MPEC is locally infeasible. If one would solve the problem to global optimality, with some $s_{i}>0$ being the optimal solution, then the MPEC is infeasible.
A special case is if the functions $c$ are concave. Then every BNLP of (26) is a convex optimization problem. If every BNLP has a solution with at least one $s_{i}>0$, then the MPEC is globally infeasible. We mention one more special case, which is hard to verify in practice. If the linearization of $c$ at $x$ is an inner approximation of $c(x)\geq 0$. Then it is sufficient to solve a single LPEC to find a feasible point of the MPEC.
3.4 Phase II: Computing a B-stationary point
Most ingredients of Phase II are already listed in Algorithm 1. A version with more implementation details is given in Algorithm 3.
Algorithm 3 Phase II of MPECopt
[…]
5:Solve $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}+d^{k+l},{\mathcal{I}}_{1}(x^{k}+d^{k+1})$ to obtain $x^{k,l}$IfBNLP solver failed or $f(x^{k,l})\geq f(x^{k})$
[…]
7:${\rho}^{k,l+1}={\gamma^{\mathrm{U}}\rho^{k,l}}$
[…]
In the next section, we prove that this algorithm terminates after a finite number of iterations. In a practical implementation, we allow a predetermined maximum number of inner and outer iterations, and if this number is not large enough to find a B-stationary point, then the solver returns an according message. A point is declared to be B-stationary if $\|d^{k,l}\|_{\infty}\leq\texttt{tol}_{\mathrm{B}}$, where $\texttt{tol}_{\mathrm{B}}=10^{-8}$. The number of inner iterations $N^{\mathrm{in}}$ must be such that $\rho^{k,l}$ cannot be less than $\texttt{tol}_{\mathrm{B}}$. In the inner iterations, the BNLP is solved again only if the active set LPEC solution predicts new active sets ${\mathcal{I}}_{1}(x^{k}+d^{k,l})$ and ${\mathcal{I}}_{2}(x^{k}+d^{k,l})$, compared to the previous LPEC solution $x^{k}+d^{k,l-1}$. A new iterate is accepted only if it strictly reduces the objective of the MPEC (1). To reduce conservatism, the radius of the trust region can be rested or increased after each successful major iteration but this is not necessary for convergence.
Recall that a B-stationary point of an MPEC is also a stationary point of all corresponding BNLPs and the TNLP, cf. Eq. (10). Therefore, instead of solving the $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}+d^{k+l},{\mathcal{I}}_{1}(x^{k}+d^{k+1})$, MPECopt would also converge if one solves a sequence of TNLPs based on the estimate ${\mathcal{I}}_{0+}(x^{k}+d^{k+l})$, ${\mathcal{I}}_{+0}(x^{k}+d^{k+l})$ and ${\mathcal{I}}_{00}(x^{k}+d^{k+l})$. In a practical implementation, we prefer to solve a sequence of BNLPs over TNLPs because of the following two reasons. First, if due to rounding errors the set ${\mathcal{I}}_{00}(x^{k}+d^{k+l})$ is computed wrongly, one can obtain infeasible subproblems. This can be avoided by estimating the BNLP index sets directly from the MILP solution, as discussed in Section 3.2.2. Second, the feasible set of all BNLPs is a superset of the TNLP, which may allow us to find better descent directions from non-stationary points.
In our implementation of Algorithm 3, we reset the trust region after every successful inner iteration and use the following constants: $\rho^{0}=10^{-3}$, $\rho^{\mathrm{lb}}=10^{-7}$, $\rho^{\mathrm{ub}}=100$, $\gamma^{\mathrm{L}}=0.1$, $\gamma^{\mathrm{U}}=10$, $N^{\mathrm{out}}=25$, and $N^{\mathrm{in}}=4$.
@@ 4 MPECopt convergence proof
This section provides the convergence proof of the MPECopt method. Before stating the main result, we state the main assumption and some auxiliary results.
@@ 4.1 Assumptions
MPECopt can use any globalized NLP solvers that satisfy the following assumption.
@@
where $f:{\mathbb{R}}^{n}\to{\mathbb{R}}$ and $g:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n_{g}}$ are twice continuously differentiable functions. Starting from an initial guess $x^{0}\in{\mathbb{R}}^{n}$, an NLP solver returns a stationary point $x^{*}$ or a certificate of local infeasibility. Moreover, if $x^{0}$ is a feasible but not stationary point of (27), then it returns a point $x^{*}$ such that $f(x^{*})<f(x^{0})$.
The NLP (27) is a generic problem that covers all NLPs solved in this paper, including (8), Reg($\tau$) and the BNLPs arising in Section 3.3.2. This assumption is met in practice by many globalized NLP solver implementations, e.g. IPOPT Waechter2006 , filterSQP Fletcher2002b , SNOPT Gill2005 and uno Vanaret2024 , to name a few. Methods that can be efficiently warm-started may profit from the fact that in Phase II all iterates are feasible.
The most restrictive requirement is that starting from a feasible point $x^{0}$, the NLP solver finds $x^{*}$ such that $f(x^{*})<f(x^{0})$. It may happen that even starting from a feasible point, the NLP solver cannot find a point acceptable to a globalization strategy, i.e., better objective and sufficiently feasible. In such cases, a feasibility restoration phase is usually triggered Waechter2006 , which may return a point with a higher objective. However, we never observed this event in any of our extensive numerical experiments, and thus we do not treat this outcome further. A possible remedy is to add the constraint $f(x)\leq f(x^{k})$ to the current BNLP.
We state our requirements on the LPEC solver.
Assumption 4.2 (LPEC solver)
Let $x^{k}\in{\mathbb{R}}^{n}$ and $\rho>0$. The LPEC solver applied to $\mathrm{LPEC}(x^{k},\rho)$ satisfies the following:
  • If $x^{k}\notin\Omega$, then it certifies that the LPEC is locally infeasible or returns a local minimizer $d\in{\mathbb{R}}^{n}$.
  • If $x^{k}\in\Omega$ and $d=0$ is a local minimizer then this is certified. Otherwise, a local optimizer $d\neq 0$ such that $\nabla f(x^{k})^{\top}d<0$ is returned.
This assumption is always satisfied by globally solving the MILP (21), but it is clear that global optimality is not necessary.
We do not treat unbounded problems, and require the following assumption, which is satisfied by adding lower and upper bounds to all variables $x\in{\mathbb{R}}^{n}$.
Assumption 4.3
In the proof below we often use Eq. (10). This relation also holds for the respective tangent cones of the MPEC, TNLP and the BNLPs. If a constraint qualification holds, this is also true for the MPEC-linearized feasible cones, which are used to define the feasible set of the LPECs. Thus, to avoid the degenerate case of Example 3.1 we require a weak constraint qualification.
Assumption 4.4
Let $x\in\Omega$ be a feasible point of the MPEC (1). Then the MPEC-Abadie CQ holds at this point.
Note that practical NLP solvers from Assumption 4.1 may need stronger CQs, e.g. the MFCQ for each BNLP, resulting in MPEC-MFCQ. However, as long as a solver can satisfy Assumption 4.1, we do not need to specify stronger assumptions for MPECopt.

### v2

@@ 3.3 Phase I: Finding a feasible branch NLP
1:Input: $x^{0}$, $\rho_{I}>0$, $\kappa\in(0,1)$, $\tau^{0}>0$, $N^{\tau}>1$
2: while do
3:   Solve $\mathrm{Reg}(\tau^{k})$ for $x^{*}(\tau^{k})$
4:   if $\mathrm{Reg}(\tau^{k})$ locally infeasible then
5:    break $\triangleright$ MPEC locally infeasible
6:   else if $x^{*}(\tau^{k})\in\Omega$ then
7:    break; return $x^{0}=x^{*}(\tau^{k})$ $\triangleright$ $\mathrm{Reg}(\tau^{k})$ solution feasible for MPEC
8:   else if $h_{\perp}(x(\tau^{k}))<\rho_{I}$ then
9:    Solve $\mathrm{LPEC}(x^{*}(\tau^{k}),\rho_{I})$ for $d^{k}$
10:    if $d^{k}$ exists then
11:      Solve $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{*}(\tau^{k})+d^{k}),{\mathcal{I}}_{2}(x^{*}(\tau^{k})+d^{k}))$ for $x^{0}$.
12:      if Solution exists then
13:       break, return $x^{0}$ $\triangleright$ feasible point found
14:   else
15:    $\tau^{k+1}=\kappa\tau^{k}$. $\triangleright$ no feasible LPEC/BNLP found, reduce homotopy parameter
We show below in Corollary 4.1 that if the MPEC admits a feasible point, then for sufficiently small $\tau$, the solution of the LPEC corresponds to the index set of a feasible BNLP. Consequently, the while loop in Algorithm 2 has finite termination. In a practical implementation, a point is declared feasible if the constraint residual is below a certain tolerance tol_h. Formally, we require $h(x)\leq\texttt{tol}_{h}$, where $h(x)$ is the total infeasibility defined as:
$\begin{split}&h(x)=\max(h_{c}(x),h_{\!\perp\!}(x)),\\ &h_{c}(x)=\max(c^{-}(x)),\ \!h_{\!\perp\!}(x)\!=\!\max(|\min(x_{1},x_{2})|),\end{split}$ (22)
with $c^{-}(x)=\min(0,c(x))$.
If $\mathrm{Reg}(\tau^{k})$ is infeasible (line 5), or its solution is feasible for the MPEC (line 7), the algorithm terminates. Otherwise, the solution $x^{*}(\tau^{k})$ is used to construct ${\mathrm{LPEC}}(x^{*}(\tau^{k}),\rho_{I})$. The fixed trust-region radius $\rho_{I}>0$ must be sufficiently large to ensure feasibility of this LPEC. Since a solution of $\mathrm{Reg}(\tau^{k})$ already satisfies the general inequality constraints, i.e., $c(x^{*}(\tau^{k}))\geq 0$, the smallest value of $\rho_{I}$ required for feasibility is $\rho_{I}>h_{\perp}(x^{*}(\tau^{k}))$.
The LPEC is solved only if this inequality holds (line 8). If it is not yet satisfied or the LPEC is infeasible, the homotopy parameter is reduced to obtain a better linearization point $x^{*}(\tau^{k})$ (line 15). If a solution to the LPEC is found, Algorithm 2 attempts to solve the corresponding $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{*}(\tau^{k})+d^{k}),{\mathcal{I}}_{2}(x^{*}(\tau^{k})+d^{k}))$ (line 11). If this is successful, a feasible point is obtained. Otherwise, the homotopy parameter is reduced to compute a better linearization point $x^{*}(\tau^{k+1})$.
Theorem 4.1 shows that if $x^{*}(\tau^{k})$ is sufficiently close to a feasible point $x\in\Omega$, then $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{*}(\tau^{k})+d^{k}),{\mathcal{I}}_{2}(x^{*}(\tau^{k})+d^{k}))$ is feasible, where $d^{k}$ is a feasible point of $\mathrm{LPEC}(x^{*}(\tau^{k}),\rho^{I})$. Moreover, under MPEC-MFCQ, the sequence $x^{*}(\tau^{k})$ converges to a stationary and thus feasible point of the MPEC Hoheisel2013 ; Kanzow2015 .
Remark 3.2
Algorithm 2 with the outcomes in lines 7 or 13 can also be viewed as a standalone method that computes a stationary point of MPEC (1c), though not necessarily a B-stationary point. However, if the computed point is B-stationary, this will be immediately verified in Phase II. If the trust-region radius is sufficiently small (cf. Lemma 3.1), the first LPEC to be solved will have a solution $d=0$. This behavior was occasionally observed in our numerical experiments in Section 5.
3.4 Phase II: Computing a B-stationary point
Algorithm 3 provides a detailed version of the Phase II algorithm outlined in Algorithm 1. It consists of two nested loops: an outer loop with counter $k$, which makes progress toward B-stationary points, and an inner loop with counter $l$, which controls the trust-region radius $\rho^{k,l}$ and solves LPECs and BNLPs until a point $x^{k,l}\in\Omega$ with improved objective value is found, i.e., $f(x^{k,l})<f(x^{k})$. Under reasonable assumptions, the inner loop is guaranteed to terminate in finitely many steps; cf. Proposition 4.3. With additional conditions, the outer loop may also terminate finitely; cf. Theorem 4.2.
Algorithm 3 Phase II of MPECopt
1:Input: $x^{0}\in\Omega$, $\gamma^{\mathrm{L}}\in(0,1)$, $N^{\mathrm{out}}$, $N^{\mathrm{in}}$, $\rho^{\mathrm{lb}}$, $\rho^{\mathrm{ub}}$.
2: for $k=0,\ldots,N^{\mathrm{out}}$ do $\triangleright$ major/outer iterations
3:   Reset trust region radius $\rho^{k,0}\in[\rho^{\mathrm{lb}},\rho^{\mathrm{ub}}]$;
4:   for $l=0,\ldots,N^{\mathrm{in}}$ do $\triangleright$ minor/inner iterations
5:    Solve $\mathrm{LPEC}(x^{k},\rho^{k,l})$ for $d^{k,l}$
6:    if $d^{k,l}=0$ then
7:      terminate $\triangleright$ B-stationary point found
8:    else
9:      Select partition $({\mathcal{I}}_{1}^{*}(x^{k}+d^{k,l}),{\mathcal{I}}_{2}^{*}(x^{k}+d^{k,l}))$ via Eq. (23)
10:      if $l=0$ or ${\mathcal{I}}_{1}^{*}(x^{k}+d^{k,l})\neq{\mathcal{I}}_{1}^{*}(x^{k}+d^{k,l-1})$ then
11:       Solve $\mathrm{BNLP}({\mathcal{I}}_{1}^{*}(x^{k}+d^{k,l}),{\mathcal{I}}_{2}^{*}(x^{k}+d^{k,l}))$ to obtain $x^{k,l}$
12:       if BNLP solver failed or $f(x^{k,l})\geq f(x^{k})$ then
13:         ${\rho}^{k,l+1}={\gamma^{\mathrm{L}}\rho^{k,l}}$; $\triangleright$ reduce TR radius
14:       else
15:         $x^{k+1}=x^{k,l}$; break inner for-loop; $\triangleright$ step accepted
16:      else
17:       ${\rho}^{k,l+1}={\gamma^{\mathrm{L}}\rho^{k,l}}$ $\triangleright$ reduce TR radius
Algorithm 3 starts from a feasible point $x^{0}\in\Omega$. At the beginning of each outer iteration, the trust-region radius is reset to $\rho^{k,0}\in[\rho^{\mathrm{lb}},\rho^{\mathrm{ub}}]$ with $\rho^{\mathrm{ub}}>\rho^{\mathrm{lb}}>0$. In theory any initial $\rho^{k,0}$ from the interval can be picked. In our implementation, we set $\rho^{k,0}=10^{-3}$, with $\rho^{\mathrm{ub}}=10^{3}$, and $\rho^{\mathrm{lb}}=10^{-7}$.
@@ 3.4 Phase II: Computing a B-stationary point
If $d^{k,l}\neq 0$, the algorithm proceeds by selecting and solving the corresponding BNLP (lines 911). If ${\mathcal{I}}_{00}(x^{k}+d^{k,l})\neq\emptyset$, the choice of the next partition is not unique. Proposition 4.3 ensures that there exists a threshold $\bar{\rho}$ such that for all $\rho^{k,l}<\bar{\rho}$, every possible partition from $\mathcal{P}(x^{k}+d^{k,l})$ yields a BNLP whose local minimizer improves the objective. Therefore, we may pick any partition from this set. We denote the selected partition by $({\mathcal{I}}_{1}^{*}(x^{k}+d^{k,l}),{\mathcal{I}}_{2}^{*}(x^{k}+d^{k,l}))\in\mathcal{P}(x^{k}+d^{k,l})$. In our implementation, we use the MILP solution to pick the partition:
${\mathcal{I}}^{*}_{1}(x^{k}+d^{k,l})=\{i\mid y_{i}^{k,l}=0\},\ {\mathcal{I}}^{*}_{2}(x^{k}+d^{k,l})=\{i\mid y_{i}^{k,l}=1\}.$ (23)
Here the integer values $y_{i}^{k,l}$ implicitly select the corresponding branches (cf. Eq. (18c)–(18d)). Alternative partitioning rules are also valid; for instance, we also implemented a rule based on the objective gradient components corresponding to complementarity variables $x_{1,i}$ and $x_{2,i}$ Kirches2022 :
$\begin{split}&\mathcal{D}_{1}^{*}(x^{k}\!+\!d^{k,l})=\{i\in{\mathcal{I}}_{00}(x^{k}\!+\!d^{k,l})\!\mid\!\nabla f_{1,i}(x^{k}\!+\!d^{k,l})\geq\nabla f_{2,i}(x^{k}+d^{k,l})\},\\ &\mathcal{D}_{2}^{*}(x^{k}+d^{k,l})={\mathcal{I}}_{00}(x^{k}+d^{k,l})\setminus\mathcal{D}_{1}^{*}(x^{k}+d^{k,l}),\\ &{\mathcal{I}}^{*}_{1}(x^{k}+d^{k,l})={\mathcal{I}}_{0+}(x^{k}+d^{k,l})\cup\mathcal{D}_{1}^{*}(x^{k}+d^{k,l})\\ &{\mathcal{I}}^{*}_{2}(x^{k}+d^{k,l})={\mathcal{I}}_{+0}(x^{k}+d^{k,l})\cup\mathcal{D}_{2}^{*}(x^{k}+d^{k,l}).\end{split}$ (24)
Since the BNLPs are in general nonconvex problems, starting from $x^{k}$ can give different objective improvements $f(x^{k})-f(x^{k,l})>0$ for different BNLPs. For convergence, it is not necessary to pick the best or a particular one, but only sufficient decrease must be guaranteed by the NLP solver, cf. Theorem 4.2. In our numerical experiments, we did not observe performance differences between using rule (23) and  (24).
If $\rho^{k,l}>\bar{\rho}$, the partition may not change between iterations, which is why line 10 checks whether a new BNLP needs to be solved. If no new partition is selected, the trust-region radius is reduced (line 17). Moreover, for larger $\rho^{k,l}$, if the current BNLP does not improve the objective (cf. Example 3.2), the trust-region radius is reduced as well (line 13). In numerical experiments, we observed that allowing larger initial trust-region radii often accelerates convergence, though they are not necessary for convergence.
@@ 4 MPECopt convergence proof
This section provides the convergence proof of the MPECopt method. Before stating the main results, we discuss the main assumption and some auxiliary results.
@@ 4.1 Assumptions
Since the feasible set of any BNLP is a subset of the MPEC’s feasible set, an unbounded MPEC can be detected via an unbounded BNLP. Most NLP solvers readily detect this. Hence, in our convergence analysis we restrict attention to bounded problems.
Assumption 4.1
Most results in the previous section required only the MPEC-ACQ. For the convergence results in this section, we require a stronger assumption.
Assumption 4.2
Then the MPEC-MFCQ holds at any feasible point of the MPEC (1).
MPEC-MFCQ implies MPEC-ACQ (Flegel2005a, , Theorem 3.1), and therefore the degenerate case of Example 3.1 is avoided. Other useful consequences are as follows.
Proposition 4.1
Suppose that Assumption 4.2 holds at $x^{*}\in\Omega$ for (1). Then:
  1. 1.
    the classical MFCQ is satisfied at $x^{*}$ by the corresponding ${\mathrm{BNLP}}({\mathcal{I}}_{1}(x^{*}),{\mathcal{I}}_{2}(x^{*}))$ for any partition $({\mathcal{I}}_{1}(x^{*}),{\mathcal{I}}_{2}(x^{*}))\in\mathcal{P}(x^{*})$,
  2. 2.
    there exists a neighborhood $\mathcal{N}$ of $x^{*}$ and a constant $\bar{\tau}>0$ such that the classical MFCQ for Reg($\tau$) is satisfied at all $x\in\mathcal{N}\cap\Omega_{\mathrm{Reg}(\tau)}$ for all $\tau\leq\bar{\tau}$, where $\Omega_{\mathrm{Reg}(\tau)}$ is the feasible set of Reg($\tau$); and,
  3. 3.
    the tangent cones of the MPEC and the BNLPs satisfy:
    $\mathcal{T}_{\Omega}(x^{*})=\bigcup_{({\mathcal{I}}_{1}^{*},{\mathcal{I}}_{2}^{*})\in\mathcal{P}^{*}}\mathcal{T}_{\Omega_{{\mathrm{BNLP}}({\mathcal{I}}_{1}^{*},{\mathcal{I}}_{2}^{*})}}(x^{*}).$ (25)
    Moreover, the left-hand side can be replaced by the MPEC-linearized feasible cone $\mathcal{F}_{\Omega}(x^{*})$, and the terms on the right-hand side by the classical linearized feasible cones of the BNLPs; cf. Eq. (40).
Proof. The first fact is (Flegel2005a, , Lemma 3.3), the second is (Hoheisel2013, , Theorem 3.2), and the third follows from (Flegel2005a, , Lemma 3.1) and (Flegel2005a, , Theorem 3.1). ∎
Many MPEC methods require solving a sequence of NLPs, e.g., Reg($\tau$) with a decreasing sequence $\{\tau^{k}\}$. In their convergence analysis, it is typically assumed that a constraint qualification holds and that a stationary point of such an NLP is available, without further specifying what assumptions on the NLP solver are required to compute such a stationary point Hoheisel2013 ; Kanzow2015 ; Ralph2004 ; Steffensen2010 . If a minimizer, and not only a stationary point, is required, then a suitable MPEC second-order sufficient condition may be assumed. Here we proceed similarly by making assumptions on the convergence of the NLP solver.
@@
where $f:{\mathbb{R}}^{n}\to{\mathbb{R}}$ and $g:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n_{g}}$ are twice continuously differentiable functions. Starting from an initial guess $x^{0}\in{\mathbb{R}}^{n}$, an NLP solver returns either:
  1. 1.
    a stationary point of (26); moreover, if $x^{0}$ is a feasible but not stationary point of (26), then it returns a point $x^{*}$ such that for some $\alpha>0$, it holds that:
    $f(x^{0})-f(x^{*})\geq\alpha>0,$ (27)
  2. 2.
    a stationary point of the feasibility problem $\min_{x}\|g^{-}(x)\|$.
This assumption is justified as follows. The NLP (26) covers all NLPs considered in this paper, including BNLPs (8) and (Reg($\tau$)), for which MFCQ holds by Assumption 4.2 and Proposition 4.1. Under the MFCQ and additional assumptions (e.g., boundedness of the Hessian, second-order sufficient conditions), convergence of stabilized sequential quadratic programming Wright2002 ; Wright1998 and interior-point methods Vicente2002 to a local minimizer is guaranteed.
The outcomes of Assumption 4.3 are met in practice by most globalized NLP solver implementations, e.g., IPOPT Waechter2006 , filterSQP (Fletcher2002b, , Theorem 7), SNOPT Gill2005 , KNITRO Byrd2006 , to name a few.
Note that the globalization strategies of all listed NLP solvers enforce sufficient decrease conditions on every step to guarantee convergence to a stationary point. When starting from a feasible point, we assume that the step towards a stationary point satisfies a sufficient decrease conditions, which we summarize in Assumption 4.3 by inequality (27).
Even when starting from a feasible point, the NLP solver may fail to produce a point satisfying the globalization strategy (i.e., sufficiently feasible with improved objective). In such cases, a feasibility restoration phase is typically triggered Waechter2006 , possibly returning a point with higher objective. We did not observe this in any of our numerical experiments and therefore do not treat it further. A possible remedy is to add the constraint $f(x)\leq f(x^{k})+\epsilon$, with $\epsilon$ small, to the current BNLP.
Lastly, we comment on the convergence of an LPEC solver. We solve the LPEC (15) via its MILP reformulation (18). MILP solvers solve LPs or certify their infeasibility, e.g., within a branch-and-bound algorithm, which always converges Wolsey1998 . Both the interior-point method and the dual-simplex method (with suitable anti-cycling rules) either solve an LP or certify its infeasibility; see, e.g., Nocedal2006 . Thus, in our setting, the LPECs can always be solved or certified infeasible. We implement this using either the commercial solver Gurobi Gurobi or the open-source solver HiGHS Huangfu2018 , both based on branch-and-bound methods.